A certain disease is caused by a recessive mutant allele. (The wild-type allele is dominant.) However, the penetrance of the disease is 75%. Two individuals known to be heterozygotes have a child. What is the probability that the child exhibits the disease?
Answer #1

Normal = AA, Aa

Affected = aa

Aa × Aa

Gametes A a
A AA normal Aa normal
a Aa normal aa affected

Probability of affected child = 1/4

But the disease is only 75% penetrant. So, the probability will be 75% of 1/4 = 75/100 × 1/4 = 3/4 × 1/4 = 3/16.

Answer is 3/16.

Please rate high.